Four years ago Sally was twice Tim's age. Four years from now, Tim will be 3/4 the age of Sally. How old is Sally now?
Answer by jorel1380(3719)

S-4=2(T-4)
T+4=3/4(S+4)
-----------
S-4=2T-8
S=2T-4
T+4=3/4(2T-4+4)
T+4=6T/4
4T+16=6T
16=2T
8=T
S=2T-4=12
Sally is 12. Tim is 8..